OpenStudy (anonymous):

which pair coordinates represent a line with a negative slope? a.(1,2)and (2,0) b.(2,3)and (4,10) c.(5,9)and(5,8)

OpenStudy (anonymous):

option a.

OpenStudy (ash2326):

slope of a line passing through two point x1,y1 and x2,y2 is given by slope=(y2-y1)/(x2-x1) let's check each option a) 1,2 and 2,0 (2-0)/(1-2)=-1 b) 2,3 and 4,10 (3-10)/(2-4)=3.5 c) 5,9 and 5,8 (9-8)/(5-5)= infinity so option A